OpenStudy (anonymous):

Find the constant term in the expansion of (2x^2 - 1/x)^6

OpenStudy (anonymous):

I'm so lost on this one, I keep seeing people talking about pascal's triangle or doing the binomial coefficient but i honestly don't understand

OpenStudy (anonymous):

@Astrophysics Hiii are you busy right now? I don't wanna bother you or anything but do you think you could help me a little with this one?

OpenStudy (anonymous):

Since the exponent is 6, you need to use the 7th row of Pascal's triangle: 1, 6, 15, 20, 15, 6, 1 Then you mulitply those coefficients by the two "inside" terms, one in increasing order from 0 to 6, and the other in decreasing order from 6 to 0. For example, the first term will be \[1(2x)^6\left( -\frac{ 1 }{ x } \right)^0\] The second will be \[6(2x)^5\left( -\frac{ 1 }{ x } \right)^1\] The third \[15(2x)^4\left( -\frac{ 1 }{ x } \right)^2\] and so on up to the 7th term. The constant term will be the one without an x

OpenStudy (anonymous):

this time we do it the easy way with no algebra if \(r=4\) and \(n-r=6-4=2\) the exponents will be the same top and bottom that will give you the constant term

OpenStudy (anonymous):

pretty much exactly what @peachpi said

OpenStudy (anonymous):

since i remember that you are writing this as \[\binom{n}{r}a^{n-r}b^r\]

OpenStudy (anonymous):

in the expansion of \[\left( x+a \right)^n\] general term \[T _{r+1}=C_{r}^{n}x ^{n-r}a^r\]

OpenStudy (anonymous):

Here \[T _{r+1}=C _{r}^{6}\left( 2 x^2 \right)^{6-r}\left( -\frac{ 1 }{ x } \right)^r\]

OpenStudy (anonymous):

Well, I "cheated" 60 Mathematica can expand the term to the 6th power in 78 micro seconds on an Apple IMac. Refer to the attachment.

OpenStudy (anonymous):

hi ok back

OpenStudy (anonymous):

@robtobey thanks, but I kinda wanted to understand it instead :p @satellite73 hey again! so nice to see you ^-^ I understand that r=4 after finding the general term. And I know you have to subtract 6 by 4, which gets you 2

OpenStudy (anonymous):

@satellite73 but the answer at the back of my book says "60". I don't understand how they found that

OpenStudy (anonymous):

@satellite73 I mean, I did binomial coefficient, so it ended up looking like the following \[\frac{ 6! }{ 4!(6!-4!) } = \]

OpenStudy (anonymous):

=15

OpenStudy (anonymous):

And so I'm confused if you're supposed to multiply that by 4, which equals 60, which is the answer OR if I'm getting the correct answer only be chance

OpenStudy (anonymous):

oh btw @peachpi if I do the binomial expansion of this really quick, will you tell me if I got it right?

OpenStudy (anonymous):

ahh hello?

zepdrix (zepdrix):

Figure it out? :)

OpenStudy (anonymous):

no ;-; pls help

zepdrix (zepdrix):

you -_-

OpenStudy (anonymous):

we can find the 60

OpenStudy (anonymous):

im sorry lol I just don't know what I'm doing. and i really don't want to do pascal's triangle.

OpenStudy (anonymous):

ok!

OpenStudy (anonymous):

first off we know \(r=4\) right

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Were you able to set up your general term? :d

zepdrix (zepdrix):

Err whatever method you were using

OpenStudy (anonymous):

yes, it's 12-3r, which is 4

OpenStudy (anonymous):

and of course \(n=6\) so on think you need to compute it \(\binom{6}{4}\)

OpenStudy (anonymous):

yeah actually i did that on a calculator but i don't want to do the calculator way

OpenStudy (anonymous):

\[\binom{6}{4}=\binom{6}{2}=\frac{6\times 5}{2}=15\]

OpenStudy (anonymous):

basically, i know that this equals 4, right? so then, like i said, i did the binomial coefficient (look above)

OpenStudy (anonymous):

yes, i did that. do you multiply 15 by 4?

OpenStudy (anonymous):

because you also have \((2x)^2=4x^2\)

zepdrix (zepdrix):

\[\large\rm (2x^2-x^{-1})^6\quad=\sum_{k=0}^{6}\left(\begin{matrix}6 \\ \rm k\end{matrix}\right)(2x^2)^{6-k}(-x^{-1})^k\]Are you forgetting about the other numbers that make up your coefficient?

OpenStudy (anonymous):

that is where the 4 comes from

zepdrix (zepdrix):

Ya, namely that fancy 2 there :d

OpenStudy (anonymous):

uh...I just got so confused lol

OpenStudy (anonymous):

ack not again !!

OpenStudy (anonymous):

@zepdrix do you plug in the 4 for the k?

OpenStudy (anonymous):

we know \(\binom{6}{4}=15\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the first term is not \(x^2\) it is \(2x^2\) right?

zepdrix (zepdrix):

\[\large\rm \left(\begin{matrix}6 \\ \rm k\end{matrix}\right)(2x^2)^{6-k}(-x^{-1})^k\]Hopefully the Binomial Theorem stuff I wrote out made sense >.< Ya you could plug in k=4 from this point. Or whatev

OpenStudy (anonymous):

you have to square all of \(2x^2\) i.e .\((2x^2)^2=4x^4\)

OpenStudy (anonymous):

that is where the 4 comes from

OpenStudy (anonymous):

@satellite73 yes, i understand where 4 is coming from, i just don't understand where 60 is coming from @zepdrix ok so like, after that, would I get 60?

OpenStudy (anonymous):

\[15\times 4=60\]

OpenStudy (anonymous):

\[\binom{6}{4}\times 4=60\] in other words

OpenStudy (anonymous):

ok, so yeah, i totally get that then, that was my question before. once you get 15, you multiply it by your r to get 60

OpenStudy (anonymous):

thank you so much satellite!!

OpenStudy (anonymous):

also thanks zepdrix too lol both of yall are amazing~

OpenStudy (anonymous):

actually it is not because you are multiplying by \(r\) it is because you have \[\binom{6}{4}(2x^2)^2\left(\frac{1}{x}\right)^4=15\times 4\]